OpenStudy (anonymous):

PLEASE HELP Verify the identity. tan x + pi/2 = -cot x

OpenStudy (mertsj):

is it tan x + pi/2 = -c0tx or tan(x+pi/2)=-cotx ?

OpenStudy (anonymous):

tan(x+pi/2)=-cotx

OpenStudy (mertsj):

\[\tan(x+\frac{\pi}{2})=\frac{\sin(x+\frac{\pi}{2})}{\cos(x+\frac{\pi}{2})}=\] \[\frac{sinxcos\frac{\pi}{2}+cosxsin \pi\frac{\pi}{2}}{cosxcos\frac{\pi}{2}-sinxsin\frac{\pi}{2}}=\frac{sinx(0)+cosx(1)}{cosx(0)-sinx(1)}=\] \[\frac{cosx}{-sinx}=-cotx\]

OpenStudy (anonymous):

thank you so much!! :D

OpenStudy (mertsj):

yw